OpenStudy (anonymous):

Are there any more functions that have a vertical asymptote of x=0 and an x-intercept at (1,0) besides the log function? @aum Can you help?

OpenStudy (aum):

Sure, it should be possible to construct any number of such functions. Vertical asymptote at x=0: 1/x has vertical asymptote at x = 0 x-intercept at (1,0). When x = 1, y = 0 when x = 1, 1/x = 1. So if we subtract 1 we will get f(x) = 0 So the function is: f(x) = 1 - 1/x or f(x) = 1/x - 1

OpenStudy (anonymous):

Ok, I understand. Thanks!